11 con in sam s

You want to obtain a sample to estimate a population proportion. Based on previous evidence, you believe the population proportion is approximately p∗=37%p∗=37%. You would like to be 99.5% confident that your esimate is within 0.5% of the true population proportion. How large of a sample size is required?

n =



Do not round mid-calculation. However, use a critical value accurate to three decimal places.
Answer #1

Solution :

Given that,

= 0.37

1 - = 1 - 0.37= 0.63

margin of error = E = 0.5% = 0.005

At 99.5% confidence level the z is ,

= 1 - 99.5% = 1 - 0.995 =0.005

/ 2 = 0.005 / 2 = 0.0025

Z/2 = Z0.0025 = 2.807

Sample size = n = (Z/2 / E)2 * * (1 - )

= (2.807 / 0.005)2 * 0.37* 0.63

= 73466

Sample size = 73466